OpenStudy (anonymous):

- 5x - 7 = -4(2x + 4) how many soloutions does the equation have. None One Two Or Infinite Soloutions

OpenStudy (anonymous):

- 5x - 7 = -4(2x + 4) -5x - 7 = -8x - 16 -5x+8x = -16 + 7 3x = -9 x = -9/3 x = -3

OpenStudy (anonymous):

get it?